Suppose the consumer has preferences of type of Leontief, where LaTeX: c\:=\:90\ell c = 90 ℓ. Also, suppose that firms make no profit, there are not taxes, and the equilibrium real wage is 10. The optimal choice of leisure (LaTeX: \ellℓ) is

a.0.9

b.9

c.1/9

d.0.1
Answer #1

C=90l

Wage =10

The optimum choice of leisure is 0.1

l+L=T

90l+10=T

l=10/90

=0.1